Hi, for the life of me I can not graph 2 periods for
y=2sec3(x-pie/6) -1
I know that sec=1/cos but on paper im not sure how to do it, nor apparently on a graphing calculator for I get the graph as a direct line going down.
My mode is in radians too so it isn't that.
I need to state the domain for that equation and I thought it would be all real numbers and that the range would be -3_< y_<1
and no idea how to determine the asymptotes. Please can you teach me? Thank you

Answer by stanbon(75887)

graph 2 periods for
y=2sec[3(x-pie/6)] -1
I know that sec=1/cos but on paper im not sure how to do it, nor apparently on a graphing calculator for I get the graph as a direct line going down.
My mode is in radians too so it isn't that.
--
Comment: You were missing a some parentheses.

----------
I need to state the domain for that equation and I thought it would be all real numbers and that the range would be -3_< y_<1
and no idea how to determine the asymptotes.
----
Domain: All Real Numbers except where cosine = 0:
3(x-(pi/6)) = 0
x = pi/6
Note: that is also where you find a vertical asymptote.
----
3(x-(pi/6)) = pi
x-(pi/6) = pi/3
x = pi/2
Another vertical asymptote at this point.
============
Cheers,
Stan H.